Can you please help me solve by using the subtitution and elimination method (unless there is no infnite number or solution.)
x + 2y = 6,
2x + y = 8

Answer by jai_kos(139)

x + 2y = 6 -->(1)
2x + y = 8-->(2)
Multiply equation(2) by 2, we get
2x + 4y = 12 -->(3)
Now subtract equation(3) with equation(2), we get
3y = 4
Now divide the above by 3, we get
y = 4/3
Now put this value in equation(2), we get
2x + (4/3) = 8
2x =8 -(4/3)
2x =20 /3
x = 20 /3 *2
x = 10/3
Hence the value of x =10/3 and y = 4/3